Service calls come to a maintenance center according to a Poisson process, and on


average, 2.7 calls are received per minute. Find the probability that


a) no more than 4 call in any minute;


b) fewer than 2 calls come in any minute;


c) more than 10 calls come in a 5-minute period.

1
Expert's answer
2022-03-25T06:37:09-0400

We have a Poisson distribution,

"\lambda=2.7;
P_t(X=k)=\cfrac{(\lambda t)^k\cdot e^{-\lambda t}}{k!}=\cfrac{(2.7t)^k\cdot e^{-2.7t}}{k!}."


a)

"P_1(X\leq4)=\\
=P_1(X=0)+P_1(X=1)+P_1(X=2)+\\
+P_1(X=3)+P_1(X=4)=\\
=\cfrac{2.7^0\cdot e^{-2.7}}{0!}+\cfrac{2.7^1\cdot e^{-2.7}}{1!}+\\
+\cfrac{2.7^2\cdot e^{-2.7}}{2!}+\cfrac{2.7^3\cdot e^{-2.7}}{3!}+\\
+\cfrac{2.7^4\cdot e^{-2.7}}{4!}=\\
=0.0672+0.1815+0.2450+0.2205+0.1488=\\=0.8629."


b)

"P_1(X<2)=\\
=P_1(X=0)+P_1(X=1)=\\
=0.0672+0.1815=0.2487."


c)

"\lambda t=2.7\cdot5=13.5;\\
 P_5(X>10)=1-P_5(X\leq10)=\\
=1-(P_5(X=0)+P_5(X=1)+...+P_5(X=10))=\\
=1-\begin{pmatrix}
 \cfrac{13.5^0}{0!}+\cfrac{13.5^1}{1!}+...\cfrac{13.5^{10}}{10!}
\end{pmatrix}\cdot e^{-13.5}=\\
=1-(1+13.5+91.13+410.06+1383.96+3736.70+\\
+8407.56+16214.59+27362.11+41043.13+\\
+55408.28)\cdot e^{-13.5}=\\
=1-0.2112=0.7888."